# Not Like Us, Hunty: Measuring Perceptions and Behavioral Effects of Minoritized Anthropomorphic Cues in LLMs

### v1

@@ Abstract.
As large language models (LLMs) increasingly adapt and personalize to diverse sets of users, there is an increased risk of systems appropriating sociolects, i.e., language styles or dialects that are associated with specific minoritized lived experiences (e.g., African American English, Queer slang). In this work, we examine whether sociolect usage by an LLM agent affects user reliance on its outputs and user perception (satisfaction, frustration, trust, and social presence). We designed and conducted user studies where 498 African American English (AAE) speakers and 487 Queer slang speakers performed a set of question-answering tasks with LLM-based suggestions in either standard American English (SAE) or their self-identified sociolect. Our findings showed that sociolect usage by LLMs influenced both reliance and perceptions, though in some surprising ways. Results suggest that both AAE and Queer slang speakers relied more on the SAE agent, and had more positive perceptions of the SAE agent. Yet, only Queer slang speakers felt more social presence from the Queer slang agent over the SAE one, whereas only AAE speakers preferred and trusted the SAE agent over the AAE one. These findings emphasize the need to test for behavioral outcomes rather than simply assume that personalization would lead to a better and safer reliance outcome. They also highlight the nuanced dynamics of minoritized language in machine interactions, underscoring the need for LLMs to be carefully designed to respect cultural and linguistic boundaries while fostering genuine user engagement and trust.
@@ 1. Introduction
With large language models (LLMs) being applied in many domains (Milanez, 2023; Lee and Yoon, 2021; Ghimire, 2024; Hancock et al., 2020; Arora et al., 2024) and interacting with more diverse users from different backgrounds (Pandey et al., 2025), there have been increased calls for them to adapt better to users and make them more human like (Wang et al., 2024c). For example, LLMs are increasingly being personalized to users’ language style (Salemi et al., 2024) and being designed to communicate in more polite ways (Ou et al., 2024). However, while this increased adaptation and human-likeness can improve user experience (Simas and Ulbricht, 2024), this causes various risks such as anthropomorphization through personalization and over-reliance (Kirk et al., 2023; Zhou et al., 2025a).
@@ 2.1. Sociolects and Biases in LLMs
AAE is a widely studied sociolect of American English spoken predominately, but not exclusively, by Black people in the United States (U.S.) (Green, 2002; Grieser, 2022; Baker-Bell, 2020). [^1] Although some refer to the language variety as African American (Vernacular) English (AAVE) or African American Language (AAL), we opt for the African American English (AAE) terminology based on previous work (Dacon, 2022; Sap et al., 2022). AAE has its own distinct grammatical and phonological features that distinguish it from SAE. Despite existing literature on AAE, it is still an underrepresented both in society as well as online spaces, which is associated with the marginalization of Black AAE speakers in predominantly White spaces (Alim et al., 2016). Biases against AAE have been widely documented in NLP systems (Sap et al., 2019; Shearer et al., 2019) and specifically LLMs (Dodge et al., 2021; Ziems et al., 2022; Xu et al., 2021; Harris et al., 2022). For instance, a recent study (Deas et al., 2023) found that current LLMs have difficulty both generating and interpreting AAE due to the lack of Black American representation in LLMs. The authors emphasized the potential harm to the AAE-speaking community and called for further development of LLMs that can effectively interact with and understand AAE speakers (Deas et al., 2023).
Queer slang, unlike AAE, is a relatively less well-documented sociolect which can be defined by words, phrases, or metaphors predominantly used by LGBTQ+ individuals in the U.S. (e.g., in queer or drag spaces; Simes, 2005; Mann, 2011). [^2] We use the term Queer slang instead of Gay slang because it encompasses more identities within the LGBTQ+ community. The emergent meaning of the word Queer to be all encompassing of LGBTQ+ individuals can be seen in recent work (Worthen, 2023). It is closer to SAE, yet could be defined as a minority sociolect due to the its ties with Queer individuals who speak it (Casey et al., 2019). Importantly, Queer slang is heavily interconnected with AAE, as Drag and Ballroom culture as well as prominent Black Queer and femme folks (e.g., RuPaul, T.S. Madison) have historically influenced mainstream Queer culture (Crutzen, 2021; Krudysz, 2023; Laing, 2021). Recent NLP work has shown documented biases in LLMs against Queer identities and Queer slang (Dodge et al., 2021; Dhingra et al., 2023). Similar to AAE, little research explores how systems interpret, generate, and classify Queer slang differently from SAE (Felkner et al., 2023).
@@ 2.2. Minority Experiences with LLMs
Along with linguistic, cultural, gender, and ability biases (Bella et al., 2024; Gadiraju et al., 2023; Lee et al., 2024) many studies have noted that AI systems and LLMs frequently fail to account for Black and LGBTQ+ identities or challenges (Dhingra et al., 2023; Ovalle et al., 2023; Mei et al., 2023; Kleiner et al., 2024; Hofmann et al., 2024). Mengesha et al. (2021) highlights significant performance disparities in AI systems’ understanding of African American language (AAL) are due to their lack of design for the Black experience. Additionally, (Ma et al., 2024) shared LLMs tend to generate overly generic or insensitive suggestions, such as advising LGBTQ+ users to come out to unsupportive parents, which risks offending and marginalizing these communities.
@@ 3. Approach: Research Questions & Variables Studied
In order to observe how sociolect use of LLMs can affect user’s conscious and subconscious perceptions of LLMs, we explored both behavioral and perceptual effects, based on previous work. The rationale for each variable will be discussed in each respective section below.
@@ 3.2. User Perception and Pairwise Preference (RQ2)
phrasing suggestion: ”adding anthropomorphic cues to the LLM’s outputs by adapting the output to be in the user’s sociolect” ? Alternatively I thought the original phrasing was good?
@@ 4.2. Experimental Procedure
Inspired by Zhou et al. (2024), during the video assessment, participants could choose to accept the LLM’s suggested responses (“Use LLM’s Response”) or rely on their own knowledge (“I’ll figure it out myself”). There was minimal framing (”You will interact with an LLM agent”) to allow participants an opportunity develop mental models based solely on the language the systems used. Participants were told they would earn points based on their ability to correctly reject incorrect LLM suggestions or rely on accurate ones. Although no points were actually calculated, this mechanism encouraged thoughtful decision-making and discouraged blind reliance on the LLM (Bansal et al., 2019; Zhou et al., 2024). The order of LLM assignment (sociolect vs. SAE) and video presentation was counterbalanced to control for order effects ( see Figure 1).
@@
Additionally, a manipulation check assessed whether participants perceived the LLM as accurately using the intended sociolect (see Appendix §G for further details). The study concluded with a post-study LLM agent preference question, where participants were asked to choose between the sociolect or SAE agent and provide a rationale or explanation for their preference.
@@ 4.3. Templated LLM Suggestions
Following the procedure in Zhou et al. (2025a), we obtained a set of SAE warmth phrases from LLMs by prompting popular LLMs to respond to factual questions with warmth prefixes (e.g., “I’d be happy to help”). We merged these LLM-generated warmth prefixes with the confidence prefixes with reliability percentages between 30% and 70% (e.g., “I would say it’s”), drawn from Zhou et al. (2024). We compiled a list of 20 warmth phrases and 14 confidence expressions (see Appendix §H), resulting in 280 unique combinations.
@@
We translated the 280 unique SAE suggestions into AAE using GPT-4’s in-context learning capabilities. This approach aligns with recent advancements in evaluating sociolect-specific text generation (Cheng et al., 2023; Cain, 2024; Jørgensen et al., 2015; Amadeus et al., 2024). For in-context learning, we sourced AAE examples from CORAAL (Kendall and Farrington, 2023), supplemented with AAE tweet pairs (Groenwold et al., 2020) and SAE-to-AAE translations (Deas et al., 2023; Kleiner et al., 2024). The detailed prompt design and implementation process for the in-context learning methodology are described in Appendix §I.
After translating the 280 warmth-epistemic combinations into AAE, we used Prolific surveys to verify authenticity, asking participants to identify which phrases aligned most closely with AAE (see Appendix §J). We presented groups of 20 phrases to batches of 5 participants at a time and participants were tasked with identifying which sociolect they felt best aligned with the translated phrases. Only phrases that were deemed to represent AAE by at least 4 out of the 5 verifiers were selected for inclusion as AAELM suggestions. This process was repeated until we achieved a diverse selection of 10 sociolect aligned phrases with unique confidence expressions (see Appendix §K for final list of AAE translations and corresponding SAE translations.)
@@
Since no existing corpora provide modern Queer slang or SAE-to-Queer slang translations, we employed persona prompting with GPT-4 to translate SAE suggestions into Queer slang. Persona prompting, where an LLM is guided to emulate the speaking style of specific individuals, has been demonstrated to effectively replicate distinctive linguistic styles (Benharrak et al., 2024; Ivey et al., 2024). For this study, we selected personas associated with well-known Queer celebrities—RuPaul, Trixie & Katya, and T.S. Madison—each known for their use of Queer slang (Crutzen, 2021; and, 2016; Kameswara and Setyawan, 2024). Using this approach, we applied the 14 SAE warmth phrases (see Table 4) to each persona output. Through persona prompting, we generated 42 unique Queer warmth phrases (see Appendix §M). This resulted in 420 unique combinations of Queer warmth phrases.
@@ 4.4. Participant Statistics and Demographics
Out of 1,007 responses, 985 participants successfully passed the sociolect screener, with 498 in the AAELM setup and 487 in the QSLM setup.We included a manipulation check to determine whether participants recognized that the LLM’s outputs were indeed in a specific sociolect. Distinguishing between those who recognized the sociolect and those who did not was critical to our analysis. Our manipulation check assessed whether participants perceived the LLM as using the sociolect successfully. This perception was essential for distinguishing between participants who bought into the sociolect attempt and those who did not. To ensure the validity of our findings, we proceeded with quantitative analysis using only the data from participants who passed the manipulation check—those who clearly perceived the intended distinction between the sociolect-speaking LLM and the SAE using LLM. However, for our qualitative analysis, we included responses from all participants, regardless of whether they passed the manipulation check, in order to capture a broader range of perceptions and preferences that were not necessarily contingent on recognizing successful sociolect usage. Participants in the AAELM group ranged from 18 to 75 years of age, with 207 identifying as men, 186 as women, and 5 as other. Participants in the QSLM group ranged from 18 to 72 years of age, with 169 identifying as men, 190 as women, and 47 as other. (See Appendix §Q).
@@ 5.1. RQ1: Does reliance on the agent depend on whether the agent speaks the user’s sociolect or not?
Our analyses showed mixed results. Among AAE-speaking participants, reliance significantly varied depending on whether the agent spoke AAE or SAE ($d=0.12$, $p<0.05$; Fig. 2(a), Table §14), however, reliance was larger on the SAELM compared to the AAELM, contradicting our hypothesis.. For Queer slang speakers, reliance showed a marginal difference depending on whether the agent used Queer slang or SAE ($d=0.08$, $p=0.09$; Fig. 2(b), Table §14), but this result did not reach statistical significance, and thus, did not support our hypothesis.
@@
For AAE speakers, the SAELM consistently was higher than the AAELM across key perception variables. AAE participants trusted the SAELM more than the AAELM ($d=0.19$, $p<0.05$; Table §16), found it more satisfying ($d=0.2$, $p<0.05$; Table §16), and experienced less frustration with it ($d=0.21$, $p<0.05$; Table §16). Interestingly, no significant difference was observed in the perceived social presence between the two agents ($d=-0.05$, $p>0.1$; Table §16).
Unlike AAE participants, Queer slang participants did not demonstrate a significant difference in trust between the Queer slang agent and the SAELM ($d=-0.03$, $p>0.1$; Table §16). However, Queer slang participants did perceive significantly greater social presence in the Queer slang agent compared to the SAELM ($d=-0.23$, $p<0.05$; Table §16). Despite this, satisfaction levels were similar between the two agents ($d=0.07$, $p>0.1$; Table §16), and participants found the SAELM significantly less frustrating than the Queer slang agent ($d=0.245$, $p<0.05$; Table §16).
Examining pairwise perception, we find that AAE participants demonstrated a clear preference for one agent over the other ($d=0.44$, $p<0.05$; Fig. 3(a), Table §15). In contrast, Queer slang participants showed no significant preference between the two agents ($d=0.009$, $p>0.1$; Fig. 3(b), Table §15).
@@ 5.3.1. Perception variables
Using Pearson correlation coefficients, we observed significant, albeit weak, correlations between reliance and perception variables for both AAE and Queer slang participants, as shown in Table 17 and Table 18. These results support our hypothesis (H3), demonstrating that perception variables such as trust, satisfaction, lack of frustration and social presence are related to reliance on the sociolect LLM. We also confirm our hypothesis that user’s have an increased reliance on the LLM agent they explicitly preferred than the reliance on the agent that participants dispreferred.
For AAE participants, trust, satisfaction, and lack of frustration significantly correlated with reliance on both the sociolect and SAE agents, with the strongest correlation observed for satisfaction ($r=0.288$, Table 17) for AAE reliance and for SAE reliance ($r=0.377$, Table 17). Social presence was also significantly correlated with reliance on the AAE agent ($r=0.188$ Table 17) but not with the SAE agent, suggesting that sociolect alignment slightly enhances the perception of social presence in this context.
For Queer slang participants, the correlations were slightly stronger overall. Trust, social presence, and satisfaction all significantly correlated with reliance on both the sociolect and SAE agents, with satisfaction showing the highest correlation for both the sociolects ($r=0.438$; Table 18) and SAE agents ($r=0.441$; Table 18). Social presence correlated more strongly with reliance on the Queer slang agent ($r=0.264$; Table 18) than with the SAE agent ($r=0.133$; Table 18), further emphasizing the role of sociolect alignment in fostering a sense of engagement.
@@
5.4. What are people’s open ended perceptions of the sociolect agent?

### v2

@@ Abstract.
As large language models (LLMs) increasingly adapt and personalize to diverse sets of users, there is an increased risk of systems appropriating sociolects, i.e., language styles or dialects that are associated with specific minoritized lived experiences (e.g., African American English, Queer slang). In this work, we examine whether sociolect usage by a LLM agent affects user reliance on its outputs and user perception (satisfaction, frustration, trust, and social presence). We designed and conducted user studies where 498 African American English (AAE) speakers and 487 Queer slang speakers performed a set of question-answering tasks with LLM-based suggestions in either standard American English (SAE) or their self-identified sociolect. Our findings showed that sociolect usage by LLMs influenced both reliance and perceptions, though in some surprising ways. Results suggest that both AAE and Queer slang speakers relied more on the SAELM, and had more positive perceptions of the SAELM. Yet, only Queer slang speakers felt more social presence from the QSLM over the SAE one, whereas only AAE speakers preferred and trusted the SAELM over the AAE one. These findings emphasize the need to test for behavioral outcomes rather than simply assume that personalization would lead to a better and safer reliance outcome. They also highlight the nuanced dynamics of minoritized language in machine interactions, underscoring the need for LLMs to be carefully designed to respect cultural and linguistic boundaries while fostering genuine user engagement and trust.
@@ 1. Introduction
With large language models (LLMs) being applied in many domains (Milanez, 2023; Lee and Yoon, 2021; Ghimire, 2024; Hancock et al., 2020; Arora et al., 2024) and interacting with more diverse users from different backgrounds (Pandey et al., 2025), there have been increased calls for them to adapt better to users and make them more human-like (Wang et al., 2024c). For example, LLMs are increasingly being personalized to users’ language style (Salemi et al., 2024) and being designed to communicate in more polite ways (Ou et al., 2024). However, while this increased adaptation and human-likeness can improve user experience (Simas and Ulbricht, 2024), this causes various risks such as anthropomorphization through personalization and over-reliance (Kirk et al., 2023; Zhou et al., 2025a).
@@ 2.1. Sociolects and Biases in LLMs
AAE is a widely studied sociolect of American English spoken predominately, but not exclusively, by Black people in the United States (U.S.) (Green, 2002; Grieser, 2022; Baker-Bell, 2020). [^1] Although some refer to the language variety as African American (Vernacular) English (AAVE) or African American Language (AAL), we opt for the African American English (AAE) terminology based on previous work (Dacon, 2022; Sap et al., 2022). AAE has its own distinct grammatical and phonological features that distinguish it from SAE. Despite existing literature on AAE, it is still an underrepresented both in society as well as online spaces, which is associated with the marginalization of Black AAE speakers in predominantly White spaces (Alim et al., 2016). Biases against AAE have been widely documented in NLP systems (Sap et al., 2019; Shearer et al., 2019) and specifically LLMs (Dodge et al., 2021; Ziems et al., 2022; Xu et al., 2021; Harris et al., 2022). For instance, a recent study by  Deas et al. (2023) found that current LLMs have difficulty both generating and interpreting AAE due to the lack of Black American representation in LLMs. The authors emphasized the potential harm to the AAE-speaking community and called for further development of LLMs that can effectively interact with and understand AAE speakers (Deas et al., 2023).
Queer slang, unlike AAE, is a relatively less well-documented sociolect which can be defined by words, phrases, or metaphors predominantly used by LGBTQ+ individuals in the U.S. (e.g., in queer or drag spaces) (Simes, 2005; Mann, 2011). [^2] We use the term Queer slang instead of Gay slang because it encompasses more identities within the LGBTQ+ community. The emergent meaning of the word Queer to be all encompassing of LGBTQ+ individuals can be seen in recent work (Worthen, 2023). It is closer to SAE, yet could be defined as a minority sociolect due to the its ties with Queer individuals who speak it (Casey et al., 2019). Importantly, Queer slang is heavily interconnected with AAE, as Drag and Ballroom culture as well as prominent Black Queer and femme folks (e.g., RuPaul, T.S. Madison) have historically influenced mainstream Queer culture (Crutzen, 2021; Krudysz, 2023; Laing, 2021). Recent NLP work has shown documented biases in LLMs against Queer identities and Queer slang (Dodge et al., 2021; Dhingra et al., 2023). Similar to AAE, little research explores how systems interpret, generate, and classify Queer slang differently from SAE (Felkner et al., 2023).
@@ 2.2. Minority Experiences with LLMs
Along with linguistic, cultural, gender, and ability biases (Bella et al., 2024; Gadiraju et al., 2023; Lee et al., 2024) many studies have noted that AI systems and LLMs frequently fail to account for Black and LGBTQ+ identities or challenges (Dhingra et al., 2023; Ovalle et al., 2023; Mei et al., 2023; Kleiner et al., 2024; Hofmann et al., 2024). Mengesha et al. (2021) highlights significant performance disparities in AI systems’ understanding of African American language (AAL) are due to their lack of design for the Black experience. Additionally, Ma et al. (2024) shared LLMs tend to generate overly generic or insensitive suggestions, such as advising LGBTQ+ users to come out to unsupportive parents, which risks offending and marginalizing these communities.
@@ 3. Approach: Research Questions & Variables Studied
In order to observe how sociolect use of LLMs can affect user’s conscious and subconscious perceptions of LLMs, we explored both behavioral and perceptual effects based on previous work. The rationale for each variable will be discussed in each respective section below.
@@ 4.2. Experimental Procedure
Inspired by Zhou et al. (2024), during the video assessment, participants could choose to accept the LLM’s suggested responses (“Use LLM’s Response”) or rely on their own knowledge (“I’ll figure it out myself”). There was minimal framing (”You will interact with a LLM agent”) to allow participants an opportunity develop mental models based solely on the language the systems used. Participants were told they would earn points based on their ability to correctly reject incorrect LLM suggestions or rely on accurate ones. Although no points were actually calculated, this mechanism encouraged thoughtful decision-making and discouraged blind reliance on the LLM (Bansal et al., 2019; Zhou et al., 2024). The order of LLM assignment (sociolect vs. SAE) and video presentation was counterbalanced to control for order effects (see Figure 1).
@@
Additionally, a manipulation check assessed whether participants perceived the LLM as accurately using the intended sociolect (see Appendix §G for further details). The study concluded with a post-study LLM agent preference question, where participants were asked to choose between the sociolect or SAELM and provide a rationale or explanation for their preference.
@@ 4.3. Templated LLM Suggestions
Following the procedure in Zhou et al. (2025a), we obtained a set of SAE warmth phrases from LLMs by prompting popular LLMs to respond to factual questions with warmth prefixes (e.g., “I’d be happy to help”). We merged these LLM-generated warmth prefixes with the confidence prefixes with reliability percentages between 30% and 70% (e.g., “I would say it’s”), drawn from Table 6 of Zhou et al. (2024). We compiled a list of 14 warmth phrases and 14 confidence expressions (see Appendix §H), resulting in 196 unique combinations.
@@
We translated the 196 unique SAE suggestions into AAE using GPT-4’s in-context learning capabilities. This approach aligns with recent advancements in evaluating sociolect-specific text generation (Cheng et al., 2023; Cain, 2024; Jørgensen et al., 2015; Amadeus et al., 2024). For in-context learning, we sourced AAE examples from CORAAL (Kendall and Farrington, 2023), supplemented with AAE tweet pairs (Groenwold et al., 2020) and SAE-to-AAE translations (Deas et al., 2023; Kleiner et al., 2024). The detailed prompt design and implementation process for the in-context learning methodology are described in Appendix §I.
After translating the 196 warmth-epistemic combinations into AAE, we used Prolific surveys to verify authenticity, asking participants to identify which phrases aligned most closely with AAE (see Appendix §J). We presented groups of 20 phrases to batches of 5 participants at a time and participants were tasked with identifying which sociolect they felt best aligned with the translated phrases. Only phrases that were deemed to represent AAE by at least 4 out of the 5 verifiers were selected for inclusion as AAELM suggestions. This process was repeated until we achieved a diverse selection of 10 sociolect aligned phrases with unique confidence expressions (see Appendix §K for final list of AAE translations and corresponding SAE translations.)
@@
Since no existing corpora provide modern Queer slang or SAE-to-Queer slang translations, we employed persona prompting with GPT-4 to translate SAE suggestions into Queer slang. Persona prompting, where a LLM is guided to emulate the speaking style of specific individuals, has been demonstrated to effectively replicate distinctive linguistic styles (Benharrak et al., 2024; Ivey et al., 2024). For this study, we selected personas associated with well-known Queer celebrities—RuPaul, Trixie & Katya, and T.S. Madison—each known for their use of Queer slang (Crutzen, 2021; and, 2016; Kameswara and Setyawan, 2024). Using this approach, we applied the 14 SAE warmth phrases (see Table 4) to each persona output. Through persona prompting, we generated 42 unique Queer warmth phrases (see Appendix §M). This resulted in 420 unique combinations of Queer warmth phrases.
@@ 4.4. Participant Statistics and Demographics
Out of 1,007 responses, 985 participants successfully passed the sociolect screener, with 498 in the AAELM setup and 487 in the QSLM setup. We included a manipulation check to determine whether participants recognized that the LLM’s outputs were indeed in a specific sociolect. Distinguishing between those who recognized the sociolect and those who did not was critical to our analysis. Our manipulation check assessed whether participants perceived the LLM as using the sociolect successfully. This perception was essential for distinguishing between participants who bought into the sociolect attempt and those who did not. To ensure the validity of our findings, we proceeded with quantitative analysis using only the data from participants who passed the manipulation check—those who clearly perceived the intended distinction between the sociolect-speaking LLM and the SAE using LLM. However, for our qualitative analysis, we included responses from all participants, regardless of whether they passed the manipulation check, in order to capture a broader range of perceptions and preferences that were not necessarily contingent on recognizing successful sociolect usage. Among those who passed the manipulation check, those in the AAELM group ranged from 18 to 75 years of age, with 207 identifying as men, 186 as women, and 3 as non-binary and 3 as other. In the QSLM group, ages ranged from 18 to 72 years of age, with 169 identifying as men, 190 as women, and 47 as other (see Appendix §Q).
@@ 5.1. RQ1: Does reliance on the agent depend on whether the agent speaks the user’s sociolect or not?
Our analyses showed mixed results. Among AAE-speaking participants, reliance significantly varied depending on whether the agent spoke AAE or SAE ($d=0.12$, $p<0.05$; Fig. 2(a), Table 14), however, reliance was larger on the SAELM compared to the AAELM, contradicting our hypothesis. For Queer slang speakers, reliance showed a marginal difference depending on whether the agent used Queer slang or SAE ($d=0.08$, $p=0.09$; Fig. 2(b), Table 14), but this result did not reach statistical significance, and thus, did not support our hypothesis.
@@
For AAE speakers, the SAELM consistently was higher than the AAELM across key perception variables. AAE participants trusted the SAELM more than the AAELM ($d=0.19$, $p<0.05$; Table 15), found it more satisfying ($d=0.2$, $p<0.05$; Table 15), and experienced less frustration with it ($d=0.21$, $p<0.05$; Table 15). Interestingly, no significant difference was observed in the perceived social presence between the two agents ($d=-0.05$, $p>0.1$; Table 15).
Unlike AAE participants, Queer slang participants did not demonstrate a significant difference in trust between the QSLM and the SAELM ($d=-0.03$, $p>0.1$; Table 15). However, Queer slang participants did perceive significantly greater social presence in the QSLM compared to the SAELM ($d=-0.23$, $p<0.05$; Table 15). Despite this, satisfaction levels were similar between the two agents ($d=0.07$, $p>0.1$; Table 15), and participants found the SAELM significantly less frustrating than the QSLM ($d=0.245$, $p<0.05$; Table 15).
Examining pairwise perception, we find that AAE participants demonstrated a clear preference for one agent over the other ($d=0.44$, $p<0.05$; Fig. 3(a), Table 16). In contrast, Queer slang participants showed no significant preference between the two agents ($d=0.009$, $p>0.1$; Fig. 3(b), Table 16).
@@ 5.3. RQ3: Are there other factors that influence user reliance?
Using Pearson correlation coefficients, we observed significant, albeit weak, correlations between reliance and perception variables for both AAE and Queer slang participants, as shown in Table 17 and Table 18. These results support our hypothesis (H3), demonstrating that perception variables such as trust, satisfaction, lack of frustration and social presence are related to reliance on the sociolect LLM. We also confirm our hypothesis that users have an increased reliance on the LLM agent they explicitly preferred than the reliance on the agent that participants dispreferred.
For AAE participants, trust, satisfaction, and lack of frustration significantly correlated with reliance on both the AAELM and SAELM, with the strongest correlation observed for satisfaction ($r=0.288$, Table 17) for AAE reliance and for SAE reliance ($r=0.377$, Table 17). Social presence was also significantly correlated with reliance on the AAELM ($r=0.188$, Table 17) but not with the SAELM, suggesting that sociolect alignment slightly enhances the perception of social presence in this context.
For Queer slang participants, the correlations were slightly stronger overall. Trust, social presence, and satisfaction all significantly correlated with reliance on both the QSLM and SAELM, with satisfaction showing the highest correlation for both the QSLM ($r=0.438$; Table 18) and SAELM ($r=0.441$; Table 18). Social presence correlated more strongly with reliance on the QSLM ($r=0.264$; Table 18) than with the SAELM ($r=0.133$; Table 18), further emphasizing the role of sociolect alignment in fostering a sense of engagement.
@@
5.4. What are people’s open-ended perceptions of the sociolect agent?